# Reinforcement Learning in Vision: A Survey

### v2

@@ Abstract
Recent advances at the intersection of reinforcement learning (RL) and visual intelligence have enabled agents that not only perceive complex visual scenes but also reason, generate, and act within them. This survey offers a critical and up-to-date synthesis of the field. We first formalize visual RL problems and trace the evolution of policy-optimization strategies from RLHF to verifiable reward paradigms, and from Proximal Policy Optimization to Group Relative Policy Optimization. We then organize more than $200$ representative works into four thematic pillars: multi-modal large language models, visual generation, unified model frameworks, and vision-language-action models. For each pillar we examine algorithmic design, reward engineering, benchmark progress, and we distill trends such as curriculum-driven training, preference-aligned diffusion, and unified reward modeling. Finally, we review evaluation protocols spanning set-level fidelity, sample-level preference, and state-level stability, and we identify open challenges that include sample efficiency, generalization, and safe deployment. Our goal is to provide researchers and practitioners with a coherent map of the rapidly expanding landscape of visual RL and to highlight promising directions for future inquiry. Resources are available at: https://github.com/weijiawu/Awesome-Visual-Reinforcement-Learning.
@@ 1 Introduction
Reinforcement Learning (RL) has achieved remarkable successes in the field of Large Language Models (LLMs) (Jaech et al., 2024; Rafailov et al., 2023), most notably through paradigms like Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) and innovative frameworks such as DeepSeek-R1 (Guo et al., 2025a). These methodologies have significantly enhanced the capabilities of LLMs, aligning generated outputs closely with human preferences and enabling nuanced, complex reasoning and interaction capabilities previously unattainable through supervised learning alone.
@@
In recent years, inspired by these remarkable achievements from LLM, there has been an explosive interest in extending the RL methodologies that proved successful for LLMs to multimodal large models, including Vision-Language Models (VLM) (Zhou et al., 2025a; Zheng et al., 2025; Zhang et al., 2025g), Vision-Language-Action models (VLA) (Lu et al., 2025b; Tan et al., 2025b; Luo et al., 2025; Yuan et al., 2025b), diffusion-based visual generation models (Fan et al., 2023b; Black et al., 2023; Zhou et al., 2025c), and unified multimodal frameworks (Mao et al., 2025; Wang et al., 2024b; 2025a), as shown in Figure 1. Multimodal models such as Gemini 2.5 (DeepMind, 2025) have leveraged RL to align visual-textual reasoning processes and produce outputs with higher semantic coherence and alignment with human judgments. Concurrently, VLA models integrating vision and language with action-oriented outputs have adopted RL to optimize complex sequential decision-making processes in interactive environments, significantly improving task-specific performance in GUI automation (Yuan et al., 2025b; Shi et al., 2025), robotic manipulation (Lu et al., 2025b), and embodied navigation (Kim et al., 2025). The rapid advancement of diffusion-based generative models has further spurred this RL-driven innovation wave. Works like ImageReward (Xu et al., 2023) have introduced reinforcement learning to enhance the semantic alignment and visual quality of generative outputs, refining diffusion-based generation through iterative feedback mechanisms derived from human preferences or automated reward critics. Moreover, unified models that blend multiple tasks, i.e., understanding, and generation, into single architectures (Mao et al., 2025; Jiang et al., 2025b) have increasingly relied on RL-driven fine-tuning, achieving generalization and task transfer previously considered challenging. Despite the substantial progress in integrating reinforcement learning with multimodal large language models, several core challenges remain. These include stabilizing policy optimization under complex reward signals, managing high-dimensional and diverse visual inputs, and designing scalable reward functions that support long-horizon decision-making. Addressing these challenges necessitates methodological innovations in both algorithmic design and evaluation protocols.
@@
In this survey, we present a comprehensive synthesis of recent advances in visual reinforcement learning within the context of multimodal large models, with a focus on the surge of research activity since 2024. We begin by revisiting foundational RL successes in language models, such as RLHF (Ouyang et al., 2022) and DeepSeek-R1 (Guo et al., 2025a), which have laid the groundwork for multimodal adaptation. Subsequently, we discuss how these strategies have evolved in the visual domain, categorizing over 200 representative works into four key domains: (i) multimodal large language models, (ii) visual generation, (iii) unified RL frameworks, and (iv) vision-language-action agents, as shown in Figure 1. Within each category, we analyze key developments in algorithmic formulations, reward modeling, and benchmarking methodologies. Finally, we identify open challenges and future directions, highlighting the need for more efficient multimodal reasoning, robust long-horizon learning strategies for VLA tasks, and scalable, high-fidelity reward signals tailored to visual generation. Through this comprehensive overview, we offer a structured overview of visual reinforcement learning to support future research and practical deployment in this rapidly evolving field.
@@ 2 Preliminary: Reinforcement Learning in LLM
Table 1: Glossary of Symbols for Visual Reinforcement Learning. It consolidates the notation that recurs across Sections §2.1§2.3, with the rightmost column pointing to each appearance of the symbol.
@@ 2.2.1 RL from Human Feedback
RLHF (Ouyang et al., 2022) extends the underlying MDP with pairwise preference data curated from human annotators. Each preference example is a triple $(p,\,y_{\mathrm{A}},\,y_{\mathrm{B}})$, where $p$ is the prompt (or state sequence) and $(y_{\mathrm{A}},y_{\mathrm{B}})$ are two candidate continuations (trajectories, images, etc.); the label $y\!\in\!\{0,1\}$ records which continuation is preferred, as shown in Figure 2.
@@
$\mathcal{L}_{\text{RM}}=-\!\!\sum_{(p,y_{\mathrm{A}},y_{\mathrm{B}})}\Bigl{[}y\,\log\sigma\!\bigl{(}R_{\phi}(p,y_{\mathrm{A}})-R_{\phi}(p,y_{\mathrm{B}})\bigr{)}+(1-y)\,\log\sigma\!\bigl{(}R_{\phi}(p,y_{\mathrm{B}})-R_{\phi}(p,y_{\mathrm{A}})\bigr{)}\Bigr{]},$ (3)
@@
Policy optimization. The policy $\pi_{\theta}$ is finally fine-tuned by maximizing (i) the learned reward, (ii) a KL penalty that keeps the policy close to the supervised-fine-tuned baseline $\pi_{\text{SFT}}$, and (iii) an optional log-likelihood regulariser on the original pre-training distribution, as introduced in InstructGPT (Ouyang et al., 2022) [^1] The coefficients $\beta$ and $\gamma$ respectively control the strength of the KL penalty and the pre-training log-likelihood term. Setting $\gamma\!=\!0$ recovers the standard PPO objective.:
$\max_{\theta}\;$ $\underbrace{\mathbb{E}_{(p,y)\sim\pi_{\theta}}\!\bigl{[}R_{\phi}(p,y)\bigr{]}}_{\text{reward}}-\beta\,\underbrace{\mathbb{E}_{p}\!\bigl{[}\mathrm{KL}\!\bigl{(}\pi_{\theta}(\cdot\!\mid\!p)\;\|\;\pi_{\text{SFT}}(\cdot\!\mid\!p)\bigr{)}\bigr{]}}_{\text{SFT\,anchoring}}+\;\gamma\,\underbrace{\mathbb{E}_{x\sim D_{\text{pretrain}}}\bigl{[}\log\pi_{\theta}(x)\bigr{]}}_{\text{pre-training log-likelihood}}\;.$ (4)
In practice, the first two terms are optimised with KL-regularised PPO over minibatches of sampled continuations, while the third term adds the pre-training gradients (“PPO-ptx” in (Ouyang et al., 2022)) to mitigate performance regressions on the original corpus.
Three-stage recipe. Most modern RLHF pipelines follow the three-stage recipe, as shown in Figure 2 (a). Step 1: Collect demonstration data, and train a supervised policy; Step 2: Collect comparison data, and train a reward model; Step 3: Optimize a policy $\pi_{\theta}$ against the reward model using PPO. The paradigm was pioneered by Christiano et al., (Christiano et al., 2017), who trained Atari and robotic agents from pairwise human preferences. Ouyang et al., (Ouyang et al., 2022) later scaled the recipe to large language models (InstructGPT) by coupling preference modeling with PPO. For vision, reward models such as ImageReward (Xu et al., 2023) and Human Preference Score (HPS) (Wu et al., 2023b) supply dense aesthetic signals that guide text-to-image diffusion and related tasks.
@@ 2.2.2 Direct Preference Optimization
Direct Preference Optimisation (DPO) (Rafailov et al., 2023) takes exactly the same pairwise-preference data as RLHF but removes the intermediate reward-model and RL loop. Instead, it derives a closed-form, supervised objective that implicitly enforces a KL constraint to a frozen reference policy $\pi_{\text{ref}}$, as shown in Figure 2 (b).
@@
$=\bigl{[}\log\pi_{\theta}(y_{\mathrm{A}}\mid p)-\log\pi_{\theta}(y_{\mathrm{B}}\mid p)\bigr{]}-\,\bigl{[}\log\pi_{\text{ref}}(y_{\mathrm{A}}\mid p)-\log\pi_{\text{ref}}(y_{\mathrm{B}}\mid p)\bigr{]}.$ (6)
@@ 2.2.3 Reinforcement Learning with Verifiable Rewards
Reinforcement Learning with Verifiable Rewards (RLVR) eliminates the subjectivity and data-collection cost of RLHF by replacing pairwise human preferences with deterministic, programmatically checkable reward signals $v:\,(p,y)\!\mapsto\!\{0,1,\dots,K\}$. Typical examples include $\tt pass/fail$}} unit tests for code synthesis, exact-match answers in mathematics, IoU/dice thresholds for segmentation, or formal output-format validators (e.g., LeetCode compiler). Because the reward is generated online by execution or metric evaluation, RLVR removes both (i) the reward-model training stage of RLHF and (ii) the contrastive surrogate loss of DPO, while still enabling substantial policy improvements beyond supervised learning (Guo et al., 2025a; Shao et al., 2024b; Lambert et al., 2024), as shown in Figure 2 (c).
@@ 2.3.1 Proximal Policy Optimization
Proximal Policy Optimization (PPO) (Schulman et al., 2017) is a first order trust region method that updates the policy $\pi_{\theta}$ while staying close to the previous policy $\pi_{\theta_{\text{old}}}$ at every step, as shown in Figure 3 (a). In text- or image-generation problems we treat a prompt $p$ as the initial state $s_{0}$ and the continuation $\{a_{1},\ldots,a_{T}\}$ as the trajectory. Define the importance-sampling ratio $\rho_{t}(\theta)$ and immediate reward $r_{\phi}$:
@@
where the KL term(latter item) keeps the updated policy $\pi_{\theta}$ from drifting too far from the frozen reference $\pi_{\text{ref}}$. $\beta$ balances exploration (via KL proximity to the frozen reference model $\pi_{\text{ref}}$) against exploitation of the reward model. A larger $\beta$ implies stricter proximity and thus safer but potentially slower learning. Then the generalized advantage estimator (GAE) (Schulman et al., 2015) produces $\hat{A}_{t}$:
@@ 2.3.2 Group Relative Policy Optimization
Group Relative Policy Optimisation (GRPO) (Shao et al., 2024b) extends PPO by discarding the learned value (critic) network and replacing it with a group-relative baseline computed from multiple outputs sampled for the same prompt. This design markedly reduces memory consumption while aligning the advantage estimator with the comparison-based reward model, as shown in Figure 3 (b).
@@
$\mathcal{L}_{\text{GRPO}}\;=$ $\mathbb{E}_{p\sim\mathcal{D}}\Bigl{[}\tfrac{1}{G}\sum_{i=1}^{G}\tfrac{1}{|a_{i}|}\sum_{t=1}^{|a_{i}|}\min\!\bigl{(}\rho_{i,t}\,\hat{A}_{i,t},\operatorname{clip}\!\bigl{(}\rho_{i,t},1-\epsilon,1+\epsilon\bigr{)}\,\hat{A}_{i,t}\bigr{)}\Bigr{]}$ (12)

### v3

# Reinforcement Learning for Large Model: A Survey

@@ Abstract
Recent advances at the intersection of reinforcement learning (RL) and visual intelligence have enabled agents that not only perceive complex visual scenes but also reason, generate, and act within them. This survey offers a critical and up-to-date synthesis of the field. We first formalize visual RL problems and trace the evolution of policy-optimization strategies from RLHF to verifiable reward paradigms, and from Proximal Policy Optimization to Group Relative Policy Optimization. We then organize more than $250$ representative works into four thematic pillars: multi-modal large language models, visual generation, unified model frameworks, and vision-language-action models. For each pillar we examine algorithmic design, reward engineering, benchmark progress, and we distill trends such as curriculum-driven training, preference-aligned diffusion, and unified reward modeling. Finally, we review evaluation protocols spanning set-level fidelity, sample-level preference, and state-level stability, and we identify open challenges that include sample efficiency, generalization, and safe deployment. Our goal is to provide researchers and practitioners with a coherent map of the rapidly expanding landscape of visual RL and to highlight promising directions for future inquiry. Resources are available at: Visual RL.
Index Terms:
Reinforcement Learning, Multimodal Model, Survey, Visual Generation.
@@ 1 Introduction
Reinforcement Learning (RL) has achieved remarkable successes in the field of Large Language Models (LLMs) [jaech2024openai, rafailov2023direct], most notably through paradigms like Reinforcement Learning from Human Feedback (RLHF) [ouyang2022training] and innovative frameworks such as DeepSeek-R1 [guo2025deepseek]. These methodologies have significantly enhanced the capabilities of LLMs, aligning generated outputs closely with human preferences and enabling nuanced, complex reasoning and interaction capabilities previously unattainable through supervised learning alone.
@@
In recent years, inspired by these remarkable achievements from LLM, there has been an explosive interest in extending the RL methodologies that proved successful for LLMs to multimodal large models, including Vision-Language Models (VLM) [zhou2025reinforced, zheng2025deepeyes, zhang2025chain], Vision-Language-Action models (VLA) [lu2025vla, tan2025interactive, luo2025gui, yuan2025enhancing], diffusion-based visual generation models [fan2023reinforcement, black2023ddpo, zhou2025dreamdpo], and unified multimodal frameworks [mao2025unirl, wang2024emu3, wang2025discrete], as shown in Figure 1. Multimodal models such as Gemini 2.5 [google_gemini2.5_2025] have leveraged RL to align visual-textual reasoning processes and produce outputs with higher semantic coherence and alignment with human judgments. Concurrently, VLA models integrating vision and language with action-oriented outputs have adopted RL to optimize complex sequential decision-making processes in interactive environments, significantly improving task-specific performance in GUI automation [yuan2025enhancing, shi2025mobilegui], robotic manipulation [lu2025vla], and embodied navigation [kim2025rapid]. The rapid advancement of diffusion-based generative models has further spurred this RL-driven innovation wave. Works like ImageReward [xu2023imagereward] have introduced reinforcement learning to enhance the semantic alignment and visual quality of generative outputs, refining diffusion-based generation through iterative feedback mechanisms derived from human preferences or automated reward critics. Moreover, unified models that blend multiple tasks, i.e., understanding, and generation, into single architectures [mao2025unirl, jiang2025co] have increasingly relied on RL-driven fine-tuning, achieving generalization and task transfer previously considered challenging. Despite the substantial progress in integrating reinforcement learning with multimodal large language models, several core challenges remain. These include stabilizing policy optimization under complex reward signals, managing high-dimensional and diverse visual inputs, and designing scalable reward functions that support long-horizon decision-making. Addressing these challenges necessitates methodological innovations in both algorithmic design and evaluation protocols.
@@
In this survey, we present a comprehensive synthesis of recent advances in visual reinforcement learning within the context of multimodal large models, with a focus on the surge of research activity since 2024. We begin by revisiting foundational RL successes in language models, such as RLHF [ouyang2022training] and DeepSeek-R1 [guo2025deepseek], which have laid the groundwork for multimodal adaptation. Subsequently, we discuss how these strategies have evolved in the visual domain, categorizing over 200 representative works into four key domains: (i) multimodal large language models, (ii) visual generation, (iii) unified RL frameworks, and (iv) vision-language-action agents, as shown in Figure 1. Within each category, we analyze key developments in algorithmic formulations, reward modeling, and benchmarking methodologies. Finally, we identify open challenges and future directions, highlighting the need for more efficient multimodal reasoning, robust long-horizon learning strategies for VLA tasks, and scalable, high-fidelity reward signals tailored to visual generation. Through this comprehensive overview, we offer a structured overview of visual reinforcement learning to support future research and practical deployment in this rapidly evolving field.
@@ 2 Preliminary: Reinforcement Learning in LLM
TABLE I: Glossary of Symbols for Visual Reinforcement Learning. It consolidates the notation that recurs across Sections §2.1§2.3, with the rightmost column pointing to each appearance of the symbol.
@@ 2.2.1 RL from Human Feedback
RLHF [ouyang2022training] extends the underlying MDP with pairwise preference data curated from human annotators. Each preference example is a triple $(p,\,y_{\mathrm{A}},\,y_{\mathrm{B}})$, where $p$ is the prompt (or state sequence) and $(y_{\mathrm{A}},y_{\mathrm{B}})$ are two candidate continuations (trajectories, images, etc.); the label $y\!\in\!\{0,1\}$ records which continuation is preferred, as shown in Figure 2.
@@
$\mathcal{L}_{\text{RM}}$ $=-\!\!\sum_{(p,y_{\mathrm{A}},y_{\mathrm{B}})}\Bigl[y\,\log\sigma\!\bigl(R_{\phi}(p,y_{\mathrm{A}})-R_{\phi}(p,y_{\mathrm{B}})\bigr)$
$\quad+(1-y)\,\log\sigma\!\bigl(R_{\phi}(p,y_{\mathrm{B}})-R_{\phi}(p,y_{\mathrm{A}})\bigr)\Bigr],$ (3)
@@
Policy optimization. The policy $\pi_{\theta}$ is finally fine-tuned by maximizing (i) the learned reward, (ii) a KL penalty that keeps the policy close to the supervised-fine-tuned baseline $\pi_{\text{SFT}}$, and (iii) an optional log-likelihood regulariser on the original pre-training distribution, as introduced in InstructGPT [ouyang2022training] [^1] The coefficients $\beta$ and $\gamma$ respectively control the strength of the KL penalty and the pre-training log-likelihood term. Setting $\gamma\!=\!0$ recovers the standard PPO objective.:
$\max_{\theta}\;$ $\underbrace{\mathbb{E}_{(p,y)\sim\pi_{\theta}}\!\bigl[R_{\phi}(p,y)\bigr]}_{\text{reward}}-\beta\,\underbrace{\mathbb{E}_{p}\!\bigl[\mathrm{KL}\!\bigl(\pi_{\theta}(\cdot\!\mid\!p)\;\|\;\pi_{\text{SFT}}(\cdot\!\mid\!p)\bigr)\bigr]}_{\text{SFT\,anchoring}}$
$\quad+\;\gamma\,\underbrace{\mathbb{E}_{x\sim D_{\text{pretrain}}}\bigl[\log\pi_{\theta}(x)\bigr]}_{\text{pre-training log-likelihood}}\;.$ (4)
In practice, the first two terms are optimised with KL-regularised PPO over minibatches of sampled continuations, while the third term adds the pre-training gradients (“PPO-ptx” in [ouyang2022training]) to mitigate performance regressions on the original corpus.
Three-stage recipe. Most modern RLHF pipelines follow the three-stage recipe, as shown in Figure 2 (a). Step 1: Collect demonstration data, and train a supervised policy; Step 2: Collect comparison data, and train a reward model; Step 3: Optimize a policy $\pi_{\theta}$ against the reward model using PPO. The paradigm was pioneered by Christiano et al., [christiano2017deep], who trained Atari and robotic agents from pairwise human preferences. Ouyang et al., [ouyang2022training] later scaled the recipe to large language models (InstructGPT) by coupling preference modeling with PPO. For vision, reward models such as ImageReward [xu2023imagereward] and Human Preference Score (HPS) [wu2023human] supply dense aesthetic signals that guide text-to-image diffusion and related tasks.
@@ 2.2.2 Direct Preference Optimization
Direct Preference Optimisation (DPO) [rafailov2023direct] takes exactly the same pairwise-preference data as RLHF but removes the intermediate reward-model and RL loop. Instead, it derives a closed-form, supervised objective that implicitly enforces a KL constraint to a frozen reference policy $\pi_{\text{ref}}$, as shown in Figure 2 (b).
@@
$=\bigl[\log\pi_{\theta}(y_{\mathrm{A}}\mid p)-\log\pi_{\theta}(y_{\mathrm{B}}\mid p)\bigr]$
$\quad-\,\bigl[\log\pi_{\text{ref}}(y_{\mathrm{A}}\mid p)-\log\pi_{\text{ref}}(y_{\mathrm{B}}\mid p)\bigr].$ (6)
@@ 2.2.3 Reinforcement Learning with Verifiable Rewards
Reinforcement Learning with Verifiable Rewards (RLVR) eliminates the subjectivity and data-collection cost of RLHF by replacing pairwise human preferences with deterministic, programmatically checkable reward signals $v:\,(p,y)\!\mapsto\!\{0,1,\dots,K\}$. Typical examples include $\tt pass/fail$}} unit tests for code synthesis, exact-match answers in mathematics, IoU/dice thresholds for segmentation, or formal output-format validators (e.g., LeetCode compiler). Because the reward is generated online by execution or metric evaluation, RLVR removes both (i) the reward-model training stage of RLHF and (ii) the contrastive surrogate loss of DPO, while still enabling substantial policy improvements beyond supervised learning [guo2025deepseek, shao2024deepseekmath, lambert2024t], as shown in Figure 2 (c).
@@ 2.3.1 Proximal Policy Optimization
Proximal Policy Optimization (PPO) [schulman2017proximal] is a first order trust region method that updates the policy $\pi_{\theta}$ while staying close to the previous policy $\pi_{\theta_{\text{old}}}$ at every step, as shown in Figure 3 (a). In text- or image-generation problems we treat a prompt $p$ as the initial state $s_{0}$ and the continuation $\{a_{1},\ldots,a_{T}\}$ as the trajectory. Define the importance-sampling ratio $\rho_{t}(\theta)$ and immediate reward $r_{\phi}$:
@@
where the KL term(latter item) keeps the updated policy $\pi_{\theta}$ from drifting too far from the frozen reference $\pi_{\text{ref}}$. $\beta$ balances exploration (via KL proximity to the frozen reference model $\pi_{\text{ref}}$) against exploitation of the reward model. A larger $\beta$ implies stricter proximity and thus safer but potentially slower learning. Then the generalized advantage estimator (GAE) [schulman2015high] produces $\hat{A}_{t}$:
@@ 2.3.2 Group Relative Policy Optimization
Group Relative Policy Optimisation (GRPO) [shao2024deepseekmath] extends PPO by discarding the learned value (critic) network and replacing it with a group-relative baseline computed from multiple outputs sampled for the same prompt. This design markedly reduces memory consumption while aligning the advantage estimator with the comparison-based reward model, as shown in Figure 3 (b).
@@
$\mathcal{L}_{\text{GRPO}}\;=$ $\mathbb{E}_{p\sim\mathcal{D}}\Bigl[\tfrac{1}{G}\sum_{i=1}^{G}\tfrac{1}{|a_{i}|}\sum_{t=1}^{|a_{i}|}\min\!\bigl(\rho_{i,t}\,\hat{A}_{i,t},$ (12)
$\operatorname{clip}\!\bigl(\rho_{i,t},1-\epsilon,1+\epsilon\bigr)\,\hat{A}_{i,t}\bigr)\Bigr]$